OpenStudy (anonymous):

25. A punter kicks a football upward with an initial speed of 48 feet per second. After how many seconds does the ball hit the ground? Use the formula h = rt − 16t2 where h represents height in feet and r represents the initial speed.

Parth (parthkohli):

Ground means height of 0, right? So, think of this like this. \( \color{Black}{\Rightarrow 0 = 48t - 16t^2 }\)

Parth (parthkohli):

Note that I have that 48 because initial speed is 48.

OpenStudy (apoorvk):

so, equate the equation of the height with zero, since the height of the ball becomes zero when it strikes the ground! and find 't' ofcourse.

OpenStudy (anonymous):

208?

Parth (parthkohli):

This is a quadratic equation.

Parth (parthkohli):

\( \color{Black}{\Rightarrow 0 = 16t(3 - t) }\) Use the zero product rule.

OpenStudy (anonymous):

then its 48t? right

OpenStudy (anonymous):

Im sorry i have no idea how to do this :(

Parth (parthkohli):

\( \color{Black}{\Rightarrow 16t = 0 \text{ or } 3 - t = 0 }\)

Parth (parthkohli):

16t = 0 t = 0 Obviously it can't take 0 seconds, so solve for 3 - t = 0

OpenStudy (anonymous):

so the answer is 3-t=0?

OpenStudy (apoorvk):

yeah. so what's 't' then?

Parth (parthkohli):

Yes, solve this further. \( \color{Black}{\Rightarrow 3 - t = 0}\) Subtract 3 from both sides then multiply both sides by -1.

OpenStudy (anonymous):

Uggg i have no idea :\

Parth (parthkohli):

\( \color{Black}{\Rightarrow -t = -3 }\) \( \color{Black}{\Rightarrow t = 3 }\)

OpenStudy (anonymous):

so its 3.................

Parth (parthkohli):

yep

OpenStudy (anonymous):

Lol thats what i got but i thought i was really wrong thanks tho! :p lol